### v1

# On the number of divisors of Mersenne numbers

Abstract.
Denote $f(n):=\sum_{1\leqslant k\leqslant n}\tau(2^{k}-1)$, where $\tau$ is the number of divisors function. Motivated by a question of Paul Erdős, we show that the sequence of ratios $f(2n)/f(n)$ is unbounded. We also present conditional results on the divergence of this sequence to infinity. Finally, we test numerically both the conjecture $f(2n)/f(n)\to\infty$ and our sufficient conditions for it to hold.
1. Introduction
The following problem is due to Paul Erdős, who posed it in the proceedings [6] of the conference Number Theory, held in Eger, Hungary in 1996. Let $\tau(n)$ denote the number of divisors of $n$ and define
$f(n):=\sum_{1\leqslant k\leqslant n}\tau(2^{k}-1)$
to be the total number of divisors of the first $n$ Mersenne numbers $12^{k}-$ 1. After expressing the feeling that $f(n)$ increases too fast to allow an asymptotic formula as $n\to\infty$, Erdős asked:
Is it true that
$f(2n)/f(n)$
tends to a limit? [6, p. 180]
Recently, the question was also posed as Problem #893 on Thomas Bloom’s website Erdős problems [3].
In this paper we will argue that $\lim_{n\to\infty}f(2n)/f(n)$ cannot be a finite number; see Corollary 2 below. We achieve this by choosing a smaller (up to a multiplicative constant) quantity $f^{\prime}(n)$, and showing that $\lim_{n\to\infty}f^{\prime}(2n)/f^{\prime}(n)=\infty$; see Theorem 1. Thus, the remaining problem is to decide whether $f(2n)/f(n)$ diverges to infinity, or it does not have any kind of limit. We will also establish two positive results in this direction, namely Theorems 3 and 4, but only conditionally on,
  • either a certain statement about the Mersenne numbers $12^{N}-$ 1 with more divisors than their predecessors (Conjecture 1),
  • or a conjectured bound on the number of prime divisors of $\Phi_{d}(2)$, where $\Phi_{d}$ is the $d$th cyclotomic polynomial (Conjecture 2).
[…]
1.1. Notation
Throughout the paper we will keep writing $\tau(n)$ for the number of divisors of $n$, while $\omega(n)$ (resp. $\Omega(n)$) will be the number of prime factors of $n$ without (resp. with) counting multiplicities. The greatest common divisor of positive integers $m$ and $n$ will be written simply as $(m,n)$, while $\phi(n)$ will stand for the number of integers $m$ such that $1\leqslant m\leqslant n$ and $(m,n)=1$ (i.e., for Euler’s totient function). Next, $\lfloor x\rfloor$ will represent the largest integer not exceeding a real number $x$ (i.e., the floor of $x$). The logarithm with base $b$ will be written as $\log_{b}$, while $\log$ will mean the natural logarithm, namely $\log_{e}$. Notation $|S|$ will be used for the number of elements of a finite set $S$. If $g$ and $h$ are two real-valued functions defined on the set of positive integers, then we write $g(n)\ll h(n)$ and $h(n)\gg g(n)$ whenever $\limsup_{n\to\infty}|g(n)/h(n)|<\infty$. Likewise, $g(n)=O(h(n))$ is yet another way of writing $g(n)\ll h(n)$, while $g(n)=o(h(n))$ simply means that $\lim_{n\to\infty}g(n)/h(n)=0$.
2. Mean value of the divisor function
We begin with possible motivation for the studied problem. Recall a trivial bound
$\frac{1}{n}\sum_{1\leqslant k\leqslant n}\tau(k)$ $=$ $\frac{1}{n}\sum_{1\leqslant\ell\leqslant n}\Big{\lfloor}\frac{n}{\ell}\Big{\rfloor}$
$=$ $\log n+O(1),$
which had been improved already by Dirichlet; see Theorem 320 in Section 18.2 of the book [8]. Because of the above asymptotics we can say that the function $\tau(n)$ behaves “on the average” as the $\log n$. If $12^{k}-$ 1 behaved like a random integer of its size with respect to the number of divisors, then one could expect that $\tau(2^{k}-1)$ behaves “on the average” like $\log(2^{k}-1)$. Then
$\sum_{1\leqslant k\leqslant n}\log(2^{k}-1)$ $=$ $\sum_{1\leqslant k\leqslant n}k\log 2+\sum_{1\leqslant k\leqslant n}\log\Big{(}1-\frac{1}{2^{k}}\Big{)}$
$=$ $\frac{n(n+1)\log 2}{2}+O\Big{(}\sum_{1\leqslant k\leqslant n}\frac{1}{2^{k}}\Big{)}$
$=$ $\frac{n^{2}\log 2}{2}+O(n)$
could lead us to further conclude that $f(2n)/f(n)$ has a finite limit, which in fact should be equal to
$\lim_{n\to\infty}\frac{(1/2)(2n)^{2}\log 2+O(n)}{(1/2)n^{2}\log 2+O(n)}=4.$
This informal reasoning makes plausible the conjecture that the ratio $f(2n)/f(n)$ stabilizes, but it is, at the same time, quite misleading. Namely, as we have already mentioned, Corollary 2 below will imply that $\lim_{n\to\infty}f(2n)/f(n)$, if it even exists, cannot be a real number. Nevertheless, the long term behaviour of $f(2n)/f(n)$ seems to be interesting and worth studying.
3. A smaller function and highly–composite numbers
There is a natural smaller function than $\tau(2^{k}-1)$ for which the analogous summatory function $f^{\prime}(n)$ has the property that $f^{\prime}(2n)/f^{\prime}(n)$ tends to infinity as $n\to\infty$. Namely, recall that Bang [2] (and also Zsigmondy [13]) proved that $12^{m}-$ 1 has a primitive prime divisor (a prime factor $p$ that does not divide $12^{\ell}-$ 1 for any positive integer $\ell<m$) for all positive integers $m$ except for $m=1,6$. Since $12^{d}-1\mid 2^{k}-$ 1 for all divisors $d$ of $k$, it follows that $12^{k}-$ 1 has at least as many prime factors as the number of divisors of $k$ different than $1$ or $6$. Thus, $\omega(2^{k}-1)\geqslant\tau(k)-2$, which leads to
$\tau(2^{k}-1)\geqslant 2^{\omega(2^{k}-1)}\geqslant\frac{1}{4}2^{\tau(k)}.$ (1)
Thus, it makes sense to consider
$f^{\prime}(n):=\sum_{1\leqslant k\leqslant n}2^{\tau(k)}$
and rather ask whether $f^{\prime}(2n)/f^{\prime}(n)$ has a limit as $n$ tends to infinity.
Theorem 1.
We have
$\lim_{n\to\infty}\frac{f^{\prime}(2n)}{f^{\prime}(n)}=\infty.$
Proof.
It suffices to show that
$\frac{\sum_{n<k\leqslant 2n}2^{\tau(k)}}{\sum_{1\leqslant k\leqslant n}2^{\tau(k)}}$
tends to infinity. Recall that a positive integer $N$ is highly–composite if $\tau(N)>\tau(m)$ for all $1\leqslant m<N$. They were studied by Ramanujan [10] and also by Erdős [5] and Nicolas [9]. Let $N$ be the largest highly–composite number not exceeding $n$. Clearly, $N>n/2$ because $\tau(2m)>\tau(m)$ holds for any positive integer $m$. Then $2N\in(n,2n]$, so
$\frac{\sum_{n<k\leqslant 2n}2^{\tau(k)}}{\sum_{1\leqslant k\leqslant n}2^{\tau(k)}}\geqslant\frac{2^{\tau(2N)}}{n2^{\tau(N)}}>\frac{2^{\tau(2N)}}{2N2^{\tau(N)}}=2^{\tau(2N)-\tau(N)+O(\log N)}.$ (2)
It is known that
$\tau(N)=2^{(1+o(1))\log N/\log\log N}\qquad{\text{\rm as }}N\to\infty.$ (3)
Namely, the upper bound holds for completely general $N$ due to a classical result by Wigert (see Theorem 317 in [8]), while the lower bound is specific to highly–composite numbers, among other alike sequences (and it easily follows from Ramanujan’s considerations in §5 of [10]). [^1] If $p_{k}$ stands for the $k$th prime and $N$ is a highly–composite number such that $2\cdot 3\cdots p_{t}\leqslant N<2\cdot 3\cdots p_{t}p_{t+1}$ for some $t\geqslant 1$, then $\tau(N)\geqslant\tau(2\cdot 3\cdots p_{t})=2^{t}=2^{\pi(p_{t})}$, where $\pi(x):=|\{p\leqslant x:p\text{ prime}\}|$ denotes the prime-counting function. On the other hand, $\log N=(1+o(1))\vartheta(p_{t})$, where $\vartheta(x):=\sum_{p\leqslant x,p\,\text{prime}}\log p$ is the Chebyshev function, so the lower bound in (3) follows from basic asymptotic properties of $\pi$ and $\vartheta$ or from Formula (25) in §5 of [10]. Writing $N$ as in §8 of [10],
$N=2^{e_{1}}3^{e_{2}}\cdots p_{t}^{e_{t}},\qquad{\text{\rm with }}e_{1}\geqslant e_{2}\geqslant\cdots\geqslant e_{t},$
where $p_{k}$ is the $k$th prime, we have
$\tau(N)=(e_{1}+1)(e_{2}+1)\cdots(e_{t}+1)\qquad{\text{\rm and}}\qquad\tau(2N)=(e_{1}+2)(e_{2}+1)\cdots(e_{t}+1).$
Thus,
$\tau(2N)-\tau(N)=(e_{2}+1)\cdots(e_{t}+1)=\frac{\tau(N)}{e_{1}+1}\gg\frac{\tau(N)}{\log N}.$
Hence, using (3) we conclude that the exponent in (2) satisfies
$\tau(2N)-\tau(N)+O(\log N)$ $\gg$ $\frac{\tau(N)}{\log N}+O(\log N)$
$=$ $\frac{2^{(1+o(1))\log N/\log\log N}}{\log N}+O(\log N)$
$=$ $2^{(1+o(1))\log N/\log\log N}$
$=$ $2^{(1+o(1))\log n/\log\log n},$
so it tends to infinity as $n\to\infty$, which is what we wanted. ∎
The following result is an immediate consequence of Theorem 1.
Corollary 2.
We have
$\limsup_{n\to\infty}\frac{f(2n)}{f(n)}=\infty,$
i.e., the sequence $(f(2n)/f(n))_{n\geqslant 1}$ is unbounded.
Proof.
If the sequence of ratios $f(2n)/f(n)$ was bounded from above by some $C<\infty$, then we would have $f(2^{m})\leqslant C^{m}$ for every integer $m\geqslant 1$. On the other hand, by (1) and Theorem 1,
$f(2^{m})\geqslant\frac{1}{4}f^{\prime}(2^{m})=\frac{1}{2}\prod_{\ell=0}^{m-1}\frac{f^{\prime}(2^{\ell+1})}{f^{\prime}(2^{\ell})}$
grows faster than $C^{m}$ as $m\to\infty$ for any finite constant $C$. ∎
4. Highly–composite Mersenne numbers and a conditional result
One may wonder about the true nature of the function $f(n)$. Is it more likely that it occasionally grows like a polynomial in $n$, as it would be the case if $\tau(2^{k}-1)$ behaved on the average like $\log(2^{k}-1)$, or is it more likely that $f(2n)/f(n)$ tends to infinity, as it was the case for the function $f^{\prime}(n)$ obtained by replacing $\tau(2^{k}-1)$ with $2^{\tau(k)}$?
Motivated by usefulness of highly–composite numbers in the proof of Theorem 1, we introduce a similar concept when the attention is restricted to Mersenne numbers only. Let us say that $N$ is an index of a highly–composite Mersenne number if $\tau(2^{N}-1)>\tau(2^{m}-1)$ holds for all $1\leqslant m<N$. Note that this definition does not mean that $12^{N}-$ 1 itself is a highly–composite number; it just means that $12^{N}-$ 1 has more divisors than any smaller Mersenne number $12^{m}-$ 1. Here is a reasonable conjecture about those numbers.
Conjecture 1.
If $N$ is restricted to indices of highly–composite Mersenne numbers, then
$\lim_{N\to\infty}\frac{\tau(2^{N}+1)}{N}=\infty.$
Some numerical evidence to support Conjecture 1 will be presented in Section 6. With this assumption we can prove the following.
Theorem 3.
If Conjecture 1 holds, then
$\lim_{n\to\infty}\frac{f(2n)}{f(n)}=\infty.$ (4)
Proof.
The proof is similar to the proof of Theorem 1. Namely, it suffices to consider
$\frac{\sum_{n<k\leqslant 2n}\tau(2^{k}-1)}{\sum_{1\leqslant k\leqslant n}\tau(2^{k}-1)}.$
Let $N$ be the largest index of a highly–composite Mersenne number that satisfies $N\leqslant n$. Clearly, $N>n/2$, because $2^{2m}-1=(2^{m}-1)(2^{m}+1)$ and so $\tau(2^{2m}-1)>\tau(2^{m}-1)$. We then have $2N\in(n,2n]$ and thus,
$\frac{\sum_{n<k\leqslant 2n}\tau(2^{k}-1)}{\sum_{1\leqslant k\leqslant n}\tau(2^{k}-1)}\geqslant\frac{\tau(2^{2N}-1)}{n\tau(2^{N}-1)}\geqslant\frac{\tau((2^{N}-1)(2^{N}+1)}{2N\tau(2^{N}-1)}=\frac{\tau(2^{N}+1)}{2N}.$
If we assume Conjecture 1, we see that the last expression diverges to $\infty$ as $n\to\infty$ (and thus also $N\to\infty$). ∎
5. Cyclotomic polynomials and another conditional result
Our next goal is to find a more concrete sufficient condition for (4). Write
$2^{n}-1=\prod_{i=1}^{t}q_{i}^{e_{i}}$
where $q_{1}<\cdots<q_{t}$ are distinct primes and $e_{1},\ldots,e_{t}$ are positive integer exponents. Clearly,
$e_{i}\leqslant\frac{\log(2^{n}-1)}{\log q_{i}}<\frac{\log 2^{n}}{\log 2}=n.$
Better bounds are known (see, for example, Theorem 1.2 in Stewart [11]), but we will not need them. We thus have that
$\tau(2^{n}-1)=\prod_{i=1}^{t}(e_{i}+1)\leqslant n^{t}=2^{t\log n/\log 2}<2^{2\omega(2^{n}-1)\log n}.$ (5)
The bottleneck of our approach consists in getting a good upper bound on $\omega(2^{n}-1)$. We write
$2^{n}-1=\prod_{d\mid n}\Phi_{d}(2),$
where
$\Phi_{d}(X)=\prod_{\begin{subarray}{c}1\leqslant h\leqslant d\\ (h,d)=1\end{subarray}}(X-e^{2\pi ih/d})\in\mathbb{Z}[X]$
is the $d$th cyclotomic polynomial. It is known that all prime factors of $\Phi_{d}(2)$ are primes $p$ such that the multiplicative order of $2$ modulo $p$ is exactly $d$ (so they are mutually disjoint for different $d$), except that $\Phi_{d}(2)$ might also be divisible by the largest prime factor of $d$. This shows
$\omega(2^{n}-1)=\sum_{d\mid n}\omega(\Phi_{d}(2))+O(\omega(n))$
and, in combination with the well-known bound [^2] We have $\frac{\log n}{\log\log n}\geqslant\frac{\log(2\cdot 3\cdots p_{\omega(n)})}{\log\log(2\cdot 3\cdots p_{\omega(n)})}=\frac{\vartheta(p_{\omega(n)})}{\log\vartheta(p_{\omega(n)})}\gg\pi(p_{\omega(n)})=\omega(n),$ where we again used an easy bound (25) from §5 of [10].
$\omega(n)\ll\frac{\log n}{\log\log n},$
implies
$\omega(2^{n}-1)=\sum_{d\mid n}\omega(\Phi_{d}(2))+O\Big{(}\frac{\log n}{\log\log n}\Big{)}.$ (6)
It remains to make an educated guess concerning $\omega(\Phi_{d}(2))$. Trivially, from the definition of $\Phi_{d}$,
$\Phi_{d}(2)\leqslant 3^{\phi(d)}$ (7)
and therefore
$\omega(\Phi_{d}(2))\ll\frac{\log 3^{\phi(d)}}{\log\log 3^{\phi(d)}}\ll\frac{\phi(d)}{\log\phi(d)},$
but the last bound is too large to be useful. Thus, we make the following conjecture.
Conjecture 2.
The inequality
$\omega(\Phi_{d}(2))\leqslant 10\log d$ (8)
holds for all positive integers $d\geqslant 2$ with at most finitely many exceptions.
Below we offer some heuristic support for this conjecture, while numerical experimentation is postponed to Section 6. Exercise 0.4 on page 12 in the book by Hall and Tenenbaum [7] shows that, uniformly in $K\geqslant 1$ and $x\geqslant 2$,
|$|\{n\leqslant x:\Omega(n)=K\}|\ll\frac{xK\log x}{2^{K}}.$
In particular,
|$|\{n\leqslant x:\Omega(n)>K\}|\ll x\log x\sum_{k>K}\frac{k}{2^{k}}\ll\frac{Kx\log x}{2^{K}}.$
We can interpret this by saying that, for a “random” positive integer $n$, the probability of having more than $K$ prime factors (even counting multiplicities) is
$\ll\frac{K\log n}{2^{K}}.$
We apply this heuristics with $n=\Phi_{d}(2)$ and $K=10\log d$, and also use (7), getting that the probability of $\omega(\Phi_{d}(2))>10\log d$ is
$\ll\frac{(10\log d)\log(\Phi_{d}(2))}{2^{10\log d}}\ll\frac{\phi(d)\log d}{d^{10\log 2}}\leqslant\frac{\log d}{d^{10\log 2-1}}\ll\frac{1}{d^{5}}.$
Since the series
$\sum_{d\geqslant 2}\frac{1}{d^{5}}$
converges, it may be indeed the case that there are only finitely many positive integers $d$ that violate (8). Therefore, Conjecture 2 is likely to hold.
We have the following theorem.
Theorem 4.
If Conjecture 2 holds, then
$\lim_{n\to\infty}\frac{f(2n)}{f(n)}=\infty.$
Proof.
We assume Conjecture 2 and suppose that $d_{0}\geqslant 2$ is an integer such that (8) holds for every integer $d\geqslant d_{0}$. It is sufficient to verify conditionally Conjecture 1 and the claim (4) will then follow from Theorem 3. Thus, suppose that $N$ is an index of a highly–composite Mersenne number.
We need to say something about $\tau(2^{N}+1)$, so let us first take a closer look at $\tau(2^{N}-1)$. By inequality (5) we have
$\tau(2^{N}-1)\leqslant 2^{2\omega(2^{N}-1)\log N},$
while, by estimates (8) and (6),
$\omega(2^{N}-1)$ $\leqslant\sum_{d<d_{0}}\omega(\Phi_{d}(2))+\sum_{\begin{subarray}{c}d\geqslant d_{0}\\ d\mid N\end{subarray}}10\log d+O\Big{(}\frac{\log N}{\log\log N}\Big{)}$
$\leqslant O(1)+10\tau(N)\log N+O(\log N)\ll\tau(N)\log N.$
Thus, under Conjecture 2 we know
$\tau(2^{N}-1)\leqslant 2^{O(\tau(N)(\log N)^{2})}.$ (9)
On the other hand, by the fact that $N$ is an index of a highly–composite Mersenne number and relation (1), we have that
$\tau(2^{N}-1)\geqslant\max_{1\leqslant k\leqslant N}\tau(2^{k}-1)\geqslant\max_{1\leqslant k\leqslant N}2^{\tau(k)-2}.$
Choose $K$ to be the largest highly–composite number not exceeding $N$. Then $K>N/2$, and by (3) we get
$\log_{2}\tau(2^{N}-1)\geqslant\tau(K)-2=2^{(1+o(1))\log K/\log\log K}=2^{(1+o(1))\log N/\log\log N}.$ (10)
Equations (9) and (10) now show that
$\tau(N)(\log N)^{2}\gg 2^{(1+o(1))\log N/\log\log N},$
so
$\tau(N)\gg 2^{(1+o(1))\log N/\log\log N}.$
Write $N=2^{a}M$, where $M$ is odd and $a\geqslant 0$ is an integer. Then we have
$\tau(N)=(a+1)\tau(M)\ll\tau(M)\log N,$
which implies
$\tau(M)\geqslant 2^{(1+o(1))\log N/\log\log N}.$
The number $N$ has $\tau(M)$ divisors $d$ such that $N/d$ is odd and $12^{d}+1\mid 2^{N}+$ 1 for all such divisors $d$. By Bang’s theorem [2] again,
$\tau(2^{N}+1)\geqslant 2^{\tau(M)-2}.$
Hence, we get that
$\frac{\tau(2^{N}+1)}{N}\geqslant\frac{2^{\tau(M)-2}}{N}=2^{\tau(M)+O(\log N)}\geqslant 2^{2^{(1+o(1))\log N/\log\log N}},$
which diverges to infinity as $N\to\infty$. ∎
6. Numerical experimentation
Figure 1. Values of $\tau(2^{n}-1)$ for $1\leqslant n\leqslant 100$.
The number of divisors of Mersenne numbers behaves quite “erratically” and exhibits occasional huge jumps; see Figure 1. In order to perform more substantial experimentation, we first computed $\tau(2^{n}-1)$ for $1\leqslant n\leqslant 250$ using the Wolfram Mathematica [12] command DivisorSigma[0,2^n-1]. For larger values of $n$ it was necessary to use faster tools for integer factorization. Dario Alpern’s integer factorization calculator Alpertron [1], which factors using elliptic curves [4, Sec. 15.10], was able to additionally compute $\tau(2^{n}-1)$ for $251\leqslant n\leqslant 390$. It enabled us to calculate the exact values of $f(2n)/f(n)$ for $1\leqslant n\leqslant 195$, which are illustrated in Figure 2. It seems that these ratios also exhibit occasional large jumps responsible for $f(2n)/f(n)\to\infty$. What we appear to be observing is that, as soon as $2n\geqslant k$ for a Mersenne number $12^{k}-$ 1 with a record-breaking number of divisors, the quotient $f(2n)/f(n)$ also becomes large, and it does not have time to decrease too much before the next huge jump of $\tau(2^{k}-1)$ occurs.
Figure 2. Numerical values of the ratios $f(2n)/f(n)$ for $1\leqslant n\leqslant 50$ (left) and $1\leqslant n\leqslant 195$ (right).
We turn to Conjecture 1. In the left half of Table 1 we made a list of all positive integers $N\leqslant 390$ that are indices of highly–composite Mersenne numbers, and also computed the corresponding ratios $\tau(2^{N}+1)/N$ in the last column. Unfortunately, interesting things start to happen when we run out of data. Thus, we also factor Mersenne numbers $12^{n}-$ 1, $n\leqslant 1000$, for which Alpertron finishes in reasonable time. [^3] The number $12^{391}-$ 1 is the first exception: Alpertron does not get past the very first step of the factorization algorithm within three hours. Counting on the fact that factorisation is fast for numbers with many (prime) divisors, which are the numbers we are looking for, we complete the list of indices $N\leqslant 1000$ of highly–composite Mersenne numbers in the right half of Table 1. The numbers $N>390$ are not $100\%$ certified to be record-breaking, but we believe that the table is quite reliable and it now readily shows the trend $\tau(2^{N}+1)/N\to\infty$. [^4] Here we ran into difficulties of not being able to factor completely $12^{780}+$ 1 and $12^{900}+$ 1, but partial factorizations by Alpertron already give large lower bounds on $\tau(2^{780}+1)$ and $\tau(2^{900}+1)$.
$N$ $\tau(2^{N}-1)$ $\tau(2^{N}+1)/N$
$1$ $1$ $2$
$2$ $2$ $1$
$4$ $4$ $0.5$
$6$ $6$ $0.6667$
$8$ $8$ $0.25$
$12$ $24$ $0.3333$
$18$ $32$ $0.8889$
$20$ $48$ $0.2$
$24$ $96$ $0.3333$
$36$ $512$ $0.4444$
$48$ $768$ $0.1667$
$60$ 4 60846084\,6084 608 $0.2667$
$72$ 8 19281928\,1928 192 $0.4444$
$84$ 9 21692169\,2169 216 $0.3810$
$810$ 2401024010\,24010 240 $0.5926$
$120$ 73 7287372873\,72873 728 $0.5333$
$144$ 262 144262144262\,144262 144 $0.8889$
$168$ 294 912294912294\,912294 912 $1.5238$
$180$ 6 291 45662914566\,291\,4566 291 456 $1.4222$
$288$ 33 554 4323355443233\,554\,43233 554 432 $0.4444$
$300$ 100 663 296100663296100\,663\,296100 663 296 $1.7067$
$360$ 1 610 612 73616106127361\,610\,612\,7361 610 612 736 $2.8444$
$N$ $\tau(2^{N}-1)$ $\tau(2^{N}+1)/N$
$420$ 57 982 058 4965798205849657\,982\,058\,49657 982 058 496 $9.7524$
$540$ 257 698 037 760257698037760257\,698\,037\,760257 698 037 760 $60.6815$
$660$ 463 856 467 968463856467968463\,856\,467\,968463 856 467 968 $397.1879$
$720$ 1 649 267 441 66416492674416641\,649\,267\,441\,6641 649 267 441 664 $45.5111$
$780$ 1 855 425 871 87218554258718721\,855\,425\,871\,8721 855 425 871 872 $\geqslant 42.0103$
$840$ 237 494 511 599 616237494511599616237\,494\,511\,599\,616237 494 511 599 616 $624.1524$
$900$ 281 474 976 710 656281474976710656281\,474\,976\,710\,656281 474 976 710 656 $\geqslant 36.4089$
Table 1. The first $22$ indices of highly–composite Mersenne numbers (left); the next $7$ likely indices of highly–composite Mersenne numbers (right).
$d$ $\omega(\Phi_{d}(2))$ $\omega(\Phi_{d}(2))/\log d$
$1$ 00
$2$ $1$ $1.4427$
$3$ $1$ $0.9102$
$4$ $1$ $0.7213$
$5$ $1$ $0.6213$
$6$ $1$ $0.5581$
$7$ $1$ $0.5139$
$8$ $1$ $0.4809$
$9$ $1$ $0.4551$
$10$ $1$ $0.4343$
$11$ $2$ $0.8341$
$12$ $1$ $0.4024$
$13$ $1$ $0.3899$
$14$ $1$ $0.3789$
$15$ $1$ $0.3693$
$16$ $1$ $0.3607$
$17$ $1$ $0.3530$
$18$ $2$ $0.6920$
$19$ $1$ $0.3396$
$20$ $2$ $0.6676$
$d$ $\omega(\Phi_{d}(2))$ $\omega(\Phi_{d}(2))/\log d$
$21$ $2$ $0.6569$
$22$ $1$ $0.3235$
$23$ $2$ $0.6379$
$24$ $1$ $0.3147$
$25$ $2$ $0.6213$
$26$ $1$ $0.3069$
$27$ $1$ $0.3034$
$28$ $2$ $0.6002$
$29$ $3$ $0.8909$
$30$ $1$ $0.2940$
$31$ $1$ $0.2912$
$32$ $1$ $0.2885$
$33$ $1$ $0.2860$
$34$ $1$ $0.2836$
$35$ $2$ $0.5625$
$36$ $2$ $0.5581$
$37$ $2$ $0.5539$
$38$ $1$ $0.2749$
$39$ $2$ $0.5459$
$40$ $1$ $0.2711$
Table 2. Numbers of prime divisors of $\Phi_{d}(2)$ for $d\leqslant 40$.
Now we turn to Conjecture 2. Table 2 presents $\omega(\Phi_{d}(2))$ and $\omega(\Phi_{d}(2))/\log d$ for $d\leqslant 40$, but this amount of data is too small to conclude anything about an upper bound for the latter quotient. Mathematica commands PrimeNu[Cyclotomic[d,2]] easily compute numbers of prime factors of $\Phi_{d}(2)$ for $d\leqslant 250$. To expand available data, we also compute $\Phi_{d}(2)$ for $251\leqslant d\leqslant 390$ in Mathematica and then factor them using Alpertron. The numbers of prime factors of all these numbers are depicted in Figure 3. It is now reasonable to suspect a logarithmic upper bound in $d$ on the numbers $\omega(\Phi_{d}(2))$. Finally, direct verification now shows that $\omega(\Phi_{d}(2))$ never exceeds $10\log d$ for $d\leqslant 390$. In fact,
$\omega(\Phi_{d}(2))\leqslant\log_{2}d\leqslant 1.45\log d\quad\text{for every }1\leqslant d\leqslant 390.$
Figure 3. Values of $\omega(\Phi_{d}(2))$ for $1\leqslant d\leqslant 390$.
Acknowledgments
V. K. was supported by the Croatian Science Foundation under the project number HRZZ-IP-2022-10-5116 (FANAP). F. L. was supported in part by the ERC Synergy Project project DynAMiCs.